### v1

Abstract
In this paper, we quantify the resilience of nonlinear dynamical systems by studying the increased energy used by all inputs of a system that suffers a partial loss of control authority, either through actuator malfunctions or through adversarial attacks. To quantify the maximal increase in energy, we introduce the notion of an energetic resilience metric. Prior work in this particular setting considers only simple linear models and not general nonlinear dynamical systems. We first characterize the mean value of the control signal in both the nominal and malfunctioning systems, which allows us to approximate the energy in the control. We then obtain a worst-case approximation of this energy for the malfunctioning system, over all malfunctioning inputs. Based on this approximation, we derive bounds on the energetic resilience metric when control authority is lost over one actuator. A simulation example on an academic nonlinear system demonstrates that the metric is useful in quantifying the resilience of the system without significant conservatism, despite the approximations used in obtaining control energies.
1 Introduction
In the process of control system design, ensuring resilience to system faults or adversarial attacks is a key issue. Control systems operate in environments with undesirable or uncontrolled influences, which often present themselves as changes to the structure of actuated inputs to the system. For instance, the disastrous crash of the F-117A Nighthawk aircraft in 1997 resulted from missing fasteners between the elevon hydraulic actuator and the wing [1], contributing to uncontrolled deflections of the wing. More recently, a software fault affected the docking of the Nauka module to the International Space Station (ISS) in 2021. An unexpected firing of the module’s thrusters led to the tilting of the ISS by up to $540^{\circ}$ for over $45$ minutes. This was described as a “loss of attitude control” by the National Aeronautics and Space Administration (NASA) [2]. To counteract this uncontrolled misfiring, other thrusters on the ISS were fired to return the ISS to its nominal position. While this issue was successfully mitigated, the mitigating action could have resulted in increased resource consumption, such as time spent and fuel consumed. Another example of an undesirable influence was discovered in a 2024 study by Lemon [3, 4] in the context of traffic light control systems suffering from potential adversarial attacks. Due to a lack of authentication software, it was found that malicious actors could gain control over the light timings in a network of traffic lights. Light timings are among the primary inputs in a large-scale traffic control system, and manipulating them can clearly lead to significant, potentially disastrous consequences.
Changes to the input structure of a control system, and in particular actuator faults, also account for $44\%$ of all faults in spacecraft attitude control systems [5, 6]. Building redundancy in actuators [7, 8] or using control reconfiguration schemes [8, 9, 6] are common strategies to address this issue, but can lead to significant costs in control system design and operation. Notably, these faults can prevent a control system from achieving a performance objective such as reachability, safety or optimality with respect to a performance criterion. Moreover, mitigating such issues with classical fault-tolerant, robust or adaptive control strategies is not guaranteed to succeed. Fault-tolerant control only addresses actuator failures limited to reduced effectiveness while still maintaining controllability [10, 11], or actuators “locked” into producing constant inputs [12]. Considering uncontrolled inputs as external disturbances within the paradigm of robust control theory may fail as shown in [13], as robust control requires disturbances to have much smaller magnitudes than the controlled inputs [14], which is not guaranteed. Fast changes in uncontrolled inputs within a short period of time also lead to the failure of adaptive control techniques [14, 15].
Motivated by the difficulties in addressing such failure modes, in this paper we consider nonlinear dynamical systems that suffer from a loss of control authority over a subset of their actuators. Such a failure separates the inputs into controlled and uncontrolled sets. Uncontrolled inputs can potentially be chosen by an adversary and take on any values in their admissible set. However, these values are measurable, and controlled inputs, which are under the authority of the system, can use these measurements to compensate for any undesirable effects from the uncontrolled inputs. In a general nonlinear system, addressing the problem of designing controlled inputs in this setting is difficult. Instead, our objective is to quantify the additional energy in the control inputs used to compensate for the effect of uncontrolled inputs, compared to the nominal case where control authority is not lost. We propose the metric of energetic resilience to quantify this additional energy.
Our line of work is closely related to the work of Bouvier and Ornik [16], who derived conditions for reachability of a target set under any uncontrolled inputs for linear systems, termed resilient reachability. In a subsequent extension [13], a control law that enables a linear system to achieve a target despite an actuator malfunction was designed using the above conditions. In [17], a quantitative resilience metric was introduced by Bouvier et al. for linear driftless systems. This metric was defined as the maximal ratio of minimal reach times to achieve a target state between the nominal and malfunctioning systems. Such a metric quantifies the maximal additional time taken by the malfunctioning system, compared to a nominal system, in achieving its objective. This metric was further used to quantify resilience in generalized integrators in [18], general linear systems in [19, 20] and cyber-physical networks in [21].
In contrast to the above works, the energetic resilience metric we use considers minimal control energies to achieve a target rather than minimal reach times as with quantitative resilience. This metric was first introduced by Padmanabhan et al. [22] to quantify the maximal cost of disturbance for a linear system affected by bounded disturbances. Subsequently, in [22], this metric was used to quantify energetic resilience for the special case of linear driftless systems.
Recent advances have also been made in addressing the issue of resilience for nonlinear systems. De Persis and Tesi [23] and Zhang et al. [24] consider Denial-of-Service attacks on control and cyber-physical systems with nonlinear dynamics. Along similar lines, Pang et al. [25] discuss resilient control for nonlinear cyber-physical systems subjected to network attacks. Zhao et al. [26] and Jia et al. [27] also address the issues of deception and Denial-of-Service attacks over cyber-physical networks. In the problem of resilience for nonlinear systems subjected to external disturbances, Saoud et al. [28] use temporal logic specifications to introduce and quantify a metric for resilience. In a similar setting, Zhou et al. [29] use robust passivity theory to design resilient controllers for nonlinear systems with bounded disturbances. We note that to the best of our knowledge, none of these recent advances address actuator malfunctions or adversaries gaining control over a set of inputs. Our objective is to address resilience in such nonlinear dynamical systems with a set of malfunctioning inputs.
The contributions of this paper are organized as follows. In Section 2, we describe the structure of the nonlinear system we analyze, as well as providing the definition of key quantities derived throughout this paper. In particular, we define a resilience metric which aims to quantify the additional energy required to compensate the effect of malfunctioning inputs. In Section 3, we obtain a characterization of the mean value of the control signal that achieves a specific task in both the nominal and malfunctioning cases. This characterization is used to approximate the minimum control energies to achieve the task in both cases, obtaining an approximate upper bound in the malfunctioning case. Further, we analyze the case when authority is lost over one actuator, obtaining the approximate malfunctioning energy in this particular setting. In Section 4, in this special case, we obtain a bound on the resilience metric which subsequently bounds from above the maximal additional energy required to achieve a ask in the malfunctioning case, compared to the nominal case. We present a simulation example of a nonlinear dynamical system in Section 5. Here, we show that the bound on the resilience metric derived in Section 4 is useful in characterizing the worst-case control energy required under a variety of malfunctioning inputs. In particular, this bound is not significantly conservative despite the approximations used in Section 3 to obtain nominal and malfunctioning energies.
1.1 Notation and Facts
The set $\mathbb{R}^{+}\coloneqq[0,\infty)$ is the set of all non-negative real numbers. For scalars $z\in\mathbb{R}$, define the $\operatorname*{sign}(\cdot)$ function as $\operatorname*{sign}(z)=z/|z|\in\{-1,+1\}$ if $z\neq 0$, with $\operatorname*{sign}(0)=0$. For vectors $z\in\mathbb{R}^{n}$, the $\operatorname*{sign}(\cdot)$ function operates elementwise on $z$. The $p$-norm of a vector $x\in\mathbb{R}^{n}$ is defined as $\|x\|_{p}\coloneqq\left(\sum_{i=1}^{n}|x_{i}|^{p}\right)^{1/p}$, with $\|x\|_{\infty}\coloneqq\max_{i}|x_{i}|$. For a matrix $L\in\mathbb{R}^{p\times q}$ with entries indexed $l_{ij}$, define the induced matrix norms $\|L\|_{1}\coloneqq\max_{j}\sum_{i=1}^{p}|l_{ij}|$ and $\|L\|_{\infty}\coloneqq\max_{i}\sum_{j=1}^{q}|l_{ij}|$. The Moore-Penrose inverse [30], also called the pseudoinverse of $L$, is denoted $L^{\dagger}$. For a piecewise continuous function $u:[0,t_{f}]\to\mathbb{R}^{p}$, its $\mathcal{L}_{2}$ norm is defined as $\|u\|_{\mathcal{L}_{2}}\coloneqq\sqrt{\int_{t=0}^{t_{f}}\|u(t)\|_{2}^{2}~{}\mathrm{d}t}$, and represents the energy in the function over the interval $[0,t_{f}]$. The mean value of such a function is defined as $\overline{u}\coloneqq\frac{1}{t_{f}}\int_{0}^{t_{f}}u(t)\mathrm{d}t$.
[…]
2 Problem Formulation
In this article, we consider continuous-time, time-invariant nonlinear dynamical systems of the form
$\dot{x}(t)=f(x(t))+g(x(t))u(t),~{}x(0)=x_{0}\neq 0,$ (1)
where the state $x(t)\in\mathbb{R}^{n}$ and the control $u(t)\in\mathbb{R}^{m+p}$ for all $t$, $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\to\mathbb{R}^{n\times(m+p)}$. We assume the functions $f$ and $g$ satisfy the following conditions:
[…]
Condition (2) is similar to a Lipschitz condition on $f$ in the $\infty$-norm. In particular, the Lipschitz constant of $f$ in the $\infty$-norm, if it exists, satisfies (2) for $D_{f}$. Condition (3) bounds $g$ everywhere in $x$. Further, the control $u$ lies in the admissible set $\operatorname*{\mathcal{U}}$, where
$\operatorname*{\mathcal{U}}\coloneqq\left\{u:\mathbb{R}^{+}\to\mathbb{R}^{m+p}:u\text{ is piecewise continuous in $t$, }\|u(t)\|_{\infty}\leq 1\text{ for all $t$}\right\}$ (4)
in line with previous works [20, 34]. The condition $\|u(t)\|_{\infty}\leq 1$ represents actuation limits in any practical control system. Conditions (2) and (3), and the admissible set $\operatorname*{\mathcal{U}}$ in (4) guarantee that a unique solution $x(t)$ to the system (1) exists [35].
Consider a malfunction that causes the system (1) to lose authority over $p$ of its $m+p$ actuators. Then, the control $u$ and the function $g$ can be split into controlled and uncontrolled components as
$\dot{x}(t)=f(x(t))+g_{c}(x(t))u_{c}(t)+g_{uc}(x(t))u_{uc}(t),$ (5)
where the subscripts $c$ and $uc$ denote controlled and uncontrolled respectively. Here, the functions $g_{c}:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ and $g_{uc}:\mathbb{R}^{n}\to\mathbb{R}^{n\times p}$ are such that $g(\cdot)=\left[g_{c}(\cdot),~{}g_{uc}(\cdot)\right]$ satisfies (3). Based on the set of admissible controls $u\in\operatorname*{\mathcal{U}}$ with $u(t)=\left[u_{c}^{\operatorname*{\mathsf{T}}}(t),~{}u_{uc}^{\operatorname*{\mathsf{T}}}(t)\right]^{\operatorname*{\mathsf{T}}}$, the controlled and uncontrolled inputs satisfy $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ and $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$, where
$\operatorname*{\mathcal{U}_{c}}$$\coloneqq\left\{u_{c}:\mathbb{R}^{+}\to\mathbb{R}^{m}:u_{c}\text{is piecewise continuous in $t$, }\|u_{c}(t)\|_{\infty}\leq 1~{}\text{for all $t$}\right\},$ (6)
$\operatorname*{\mathcal{U}_{uc}}$$\coloneqq\left\{u_{uc}:\mathbb{R}^{+}\to\mathbb{R}^{p}:u_{uc}\text{ is piecewise continuous in $t$, }\|u_{uc}(t)\|_{\infty}\leq 1~{}\text{for all $t$}\right\}.$ (7)
In this setting, we consider the task of finite-time reachability, i.e., achieving a specified target state $x(t_{f})=x_{tg}\in\mathbb{R}^{n}$ from an initial state $x(0)=x_{0}\in\mathbb{R}^{n}$ at a specified final time $t_{f}$. Under the nominal dynamics (1), the system uses the input $u\in\operatorname*{\mathcal{U}}$, while under the malfunctioning dynamics (5), the system must use only the controlled input $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ to achieve this task. We note that in the latter case, the uncontrolled input $u_{uc}$ is arbitrarily chosen — potentially by an adversary — in the space $\operatorname*{\mathcal{U}_{uc}}$. However, we assume this choice of $u_{uc}$ is observable, potentially a priori, and hence its effect can be compensated for in the design of $u_{c}$. Achieving this task despite the effect of $u_{uc}$ corresponds to the notion of resilient reachability, formally defined below.
Definition 1.
A target $x_{tg}\in\mathbb{R}^{n}$ is resiliently reachable at a given final time $t_{f}$ from $x_{0}\in\mathbb{R}^{n}$ if for all uncontrolled inputs $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$ there exists a controlled input $u_{c}\in\operatorname*{\mathcal{U}_{c}}$, possibly dependent on $u_{uc}$, such that $x(t_{f})=x_{tg}$.
Then, our first objective can be stated as follows.
Problem 1.
Characterize the control input $u\in\operatorname*{\mathcal{U}}$ in (1) (resp. $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ in (5)) that ensures a target $x_{tg}\in\mathbb{R}^{n}$ is reachable (resp. resiliently reachable).
In Section 3, we use the solution of the nominal (1) and malfunctioning (5) systems to obtain a condition on the controls $u\in\operatorname*{\mathcal{U}}$ and $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ that achieve finite-time reachability. Further, in both the nominal and malfunctioning cases, we are interested in control signals that use minimum energy to achieve $x(t_{f})=x_{tg}$. In such a context, we make the following definitions, also used in [34].
Definition 2 (Nominal Energy).
The nominal energy is the minimum energy in the control input $u\in\operatorname*{\mathcal{U}}$ such that a target $x_{tg}$ is reached from $x_{0}$ in a given time $t_{f}$ in the system (1):
[…]
Definition 3 (Malfunctioning Energy).
The malfunctioning energy is the minimum energy in the controlled input $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ such that a target $x_{tg}$ is reached from $x_{0}$ in a given time $t_{f}$ and for a given uncontrolled input $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$, in the system (5):
[…]
the dependence of $u_{c}$ on $u_{uc}$ is explicitly denoted with a slight abuse of notation.
In (9), the controlled input $u_{c}$ may depend not only on past and current values of $u_{uc}$, but potentially also future values of this uncontrolled input. The dependence on future values considers any prior knowledge we may have about uncontrolled inputs, such as the corresponding actuators locking in place or acting with reduced effectiveness [11, 12]. The controlled input $u_{c}$ seeks to compensate for the effect of $u_{uc}$ and ensure that $x_{tg}$ is reachable from $x_{0}$ in time $t_{f}$, while also using minimal energy (9). On the other hand, the uncontrolled input $u_{uc}$ also contributes to the energy used by the actuators of the system. In a system (5) that has lost authority over some of its actuators, it is important to characterize the maximal additional energy that may be used by all actuators. In practical cases, this maximal additional energy is directly related to consumption of resources such as fuel in vehicular systems. Quantifying such an increase in resource consumption is essential to understand the maximum capacity of resources required. In an adversarial setting, the uncontrolled input $u_{uc}$ may be chosen to maximize the total energy used by the system. We thus use the following definitions, first introduced in [34], that consider the total energy used by all actuators of the system (5).
Definition 4 (Total Energy).
The total energy is the energy used by all actuators of the system (5) for a given uncontrolled input $u_{uc}$, assuming the optimal controlled input $u_{c}(u_{uc})$ in (9) is used:
$E_{T}(x_{0},x_{tg},t_{f},u_{uc})\coloneqq E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})+\|u_{uc}\|_{\mathcal{L}_{2}}^{2}.$ (10)
Definition 5 (Worst-case Total Energy).
The worst-case total energy is the maximal value of the total energy (10) over all possible uncontrolled inputs $u_{uc}$:
$\overline{E}_{T}(x_{0},x_{tg},t_{f})\coloneqq\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}E_{T}(x_{0},x_{tg},t_{f},u_{uc})=\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}\left\{E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})+\|u_{uc}\|_{\mathcal{L}_{2}}^{2}\right\}.$ (11)
The worst-case total energy quantifies the maximal effect of $u_{uc}$ on the total energy used by all actuators of the system (5).
Finally, note that the worst-case total energy (11) can be considerably larger than the nominal energy (8) due to the action of the uncontrolled input $u_{uc}$. We thus arrive at our second objective.
Problem 2.
Quantify the maximal increase in energy used by the malfunctioning system (5), in comparison to the nominal system (1).
To quantify such an increase in energy, we define the following energetic resilience metric, similar to that defined in [22, 34].
Definition 6 (Energetic Resilience).
The energetic resilience of system (5) is defined for an initial condition $x_{0}$ no farther than a distance of $R$ from the target state $x_{tg}$, as
$r_{A}(t_{f},R)\coloneqq\sup_{\|x_{0}-x_{tg}\|_{2}\leq R}\overline{E}_{T}(x_{0},x_{tg},t_{f})-E_{N}^{*}(x_{0},x_{tg},t_{f}).$ (12)
The significance of this metric is evident. Based on the definition (12), the malfunctioning system (5) will use at most $r_{A}(t_{f},R)$ more energy than the nominal system (1) to achieve finite-time reachability, for a given final time $t_{f}$ and $x_{0}$ no farther than $R$ from $x_{tg}$. The constraint $\|x_{0}-x_{tg}\|_{2}\leq R$ is required to ensure that the value of $r_{A}(t_{f},R)$ is finite.
Remark 1.
In the nonlinear systems (1) and (5), deriving a closed-form expression for the control that achieves $x(t_{f})=x_{tg}$ is generally intractable (see, e.g., [35, 36]). Similarly, closed-form expressions for the minimum energies (8) and (9) and subsequently the total (10) and worst-case total (11) energies are generally not derivable [37]. To address this, we derive a condition on the mean value of the control in both the nominal and malfunctioning cases in Section 3. This condition achieves our first objective in Problem 1, and is used to approximate the minimum energies and worst-case total energy, and these approximations are used to derive a bound on the energetic resilience metric in Section 4.
3 Approximations for Nominal, Malfunctioning and Worst-case Total Energies
In this section, we derive approximations for the nominal, malfunctioning and worst-case total energies defined in Section 2. This approximation uses a characterization of the mean value of the control signal in the interval $[0,t_{f}]$, which is derived from the system equations (1) and (5).
3.1 Nominal and Malfunctioning Energies
First consider the nominal system (1). The state trajectory of this system satisfies
$x(t)=x_{0}+\int_{0}^{t}f(x(\tau))\mathrm{d}\tau+\int_{0}^{t}g(x(\tau))u(\tau)\mathrm{d}\tau.$ (13)
Let $\tilde{x}\coloneqq x_{0}-x_{tg}$, $B=g(x_{0})$ and $f_{0}=f(x_{0})$. If $x(t_{f})=x_{tg}$,
$\tilde{x}+t_{f}B\overline{u}=-\int_{0}^{t_{f}}f(x(t))\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t,$ (14)
where we use the property (3) for $g$, and $\overline{u}=\frac{1}{t_{f}}\int_{0}^{t_{f}}u(t)\mathrm{d}t$ denotes the mean value of $u$ in the interval $[0,t_{f}]$. Let $v\equiv v(t_{f},u)\coloneqq-\int_{0}^{t_{f}}f(x(t))\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t$. If $v$ was a fixed quantity with respect to $u$ and hence $\overline{u}$, (14) is linear in $\overline{u}$ and we can write the following least-norm solution [38]:
$\overline{u}^{LS}=\frac{1}{t_{f}}B^{\dagger}(v-\tilde{x}),$ (15)
where $B^{\dagger}$ is the pseudoinverse of $B$, defined as in [30]. Note that such a least-norm solution is guaranteed to satisfy (14) only if $v-\tilde{x}$ lies in the range space of $B$, and we assume this is true. Next, the solution (15) assumes $v$ is a fixed quantity with respect to $\overline{u}$, which is clearly not true as $v$ depends directly on $u$. Nevertheless, we use the norm of $\overline{u}^{LS}$ as defined in (15) to approximate the nominal energy later in this section. A key step towards such an approximation is to show that $v$ is bounded, which is proved in the following proposition.
Proposition 1.
The quantity $v=-\int_{0}^{t_{f}}f(x(t))\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t$ is bounded as
$\|v\|_{\infty}\leq\overline{v}\coloneqq\frac{1}{D_{f}}\left[c(e^{t_{f}D_{f}}-1)-t_{f}D_{f}\|B\|_{\infty}\right]$ (16)
where $c=\|f_{0}\|_{\infty}+\|B\|_{\infty}+D_{g}$.
Proof.
See Appendix A. ∎
We now use the bound $\overline{v}$ to analyze conditions under which the constraint $u\in\operatorname*{\mathcal{U}}$ is satisfied. Note that $u\in\operatorname*{\mathcal{U}}$ implies $\|\overline{u}\|_{\infty}\leq 1$. For a fixed $v$, $\overline{u}^{LS}$ as defined in (15) provides an expression for $\overline{u}$ and we can check the condition $\|\overline{u}^{LS}\|_{\infty}\leq 1$. To consider all possible $v$, we use the bound $\|v\|_{\infty}\leq\overline{v}$ and check the following condition for a given $t_{f}$:
$\max_{\|v\|_{\infty}\leq\overline{v}}\frac{1}{t_{f}}\left\|B^{\dagger}(v-\tilde{x})\right\|_{\infty}\leq 1.$ (17)
If (17) is satisfied, then $\|\overline{u}^{LS}\|_{\infty}\leq 1$ is true for all $v(t_{f},u)$, which is a necessary condition to ensure $u\in\operatorname*{\mathcal{U}}$ for a given $t_{f}$. Furthermore, (17) is equivalent to checking
$\max_{\|v\|_{\infty}\leq\overline{v}}\frac{1}{t_{f}}\left[B^{\dagger}(v-\tilde{x})\right]_{i}\leq 1$
and $\max_{\|v\|_{\infty}\leq\overline{v}}-\frac{1}{t_{f}}\left[B^{\dagger}(v-\tilde{x})\right]_{i}\leq 1$
for each $i=1,\ldots,m+p$, where $[\cdot]_{i}$ denotes the $i$-th component of the argument. Each of the above conditions is linear in $v$, and thus can be checked efficiently on the vertices of the hypercube $\|v\|_{\infty}\leq\overline{v}$ [39].
We now use (15) to obtain the nominal energy (8). For a general nonlinear system of the form (1), obtaining closed-form expressions for optimal control signals is generally intractable [37]. However, the norm of $\overline{u}^{LS}$ for a given $v$ can be used to approximate the nominal energy using the following lemma.
Lemma 1.
For a piecewise continuous function $z:[0,t_{f}]\to\mathbb{R}^{n_{z}}$ with a given mean value $\overline{z}\in\mathbb{R}^{n_{z}}$, the energy of $z$ is related to its mean value by
$\int_{0}^{t_{f}}z^{T}(t)z(t)\mathrm{d}t\geq t_{f}\|\overline{z}\|_{2}^{2},$ (18)
where equality holds when $z(t)=\overline{z}$ for all $t\in[0,t_{f}]$.
Proof.
Since the squared norm $\|\cdot\|_{2}^{2}$ is a convex function, we have $\|\overline{z}\|_{2}^{2}=\left\|\frac{1}{t_{f}}\int_{0}^{t_{f}}z(t)\mathrm{d}t\right\|_{2}^{2}\leq\frac{1}{t_{f}}\int_{0}^{t_{f}}\|z(t)\|_{2}^{2}\mathrm{d}t,$ using Jensen’s inequality [32]. From this, the result follows. ∎
Any control signal $u$ that achieves finite-time reachability must satisfy (14). Subsequently, its mean value $\overline{u}$ must satisfy (15). Thus, the optimal control signal with minimum energy $E_{N}^{*}(x_{0},x_{tg},t_{f})$ that achieves finite-time reachability must have mean value $\overline{u}^{LS}$, where we recall that $\|v\|_{\infty}\leq\overline{v}$. Using Lemma 1, $E_{N}^{*}(x_{0},x_{tg},t_{f})\geq t_{f}\|\overline{u}^{LS}\|_{2}^{2}$. We now make the central approximation in this article, namely, that
$E_{N}^{*}(x_{0},x_{tg},t_{f})\approx t_{f}\|\overline{u}^{LS}\|_{2}^{2}.$ (19)
Note that this approximation is exact when the optimal control is constant, i.e., $u^{*}(t)=\overline{u}^{LS}$ for all $t\in[0,t_{f}]$. In particular, this approximation is exact in the case of linear, driftless systems where $D_{f}=D_{g}=0$ and $f_{0}=0$, as shown in [34] where the optimal control $u^{*}(t)=\overline{u}_{LS}$ for all $t\in[0,t_{f}]$. We also use such an approximation later in the malfunctioning case. These approximations are used to circumvent the difficulties in obtaining closed-form expressions for optimal control signals and the corresponding minimum energies [37]. Under the assumption that this approximation is exact, we obtain
$E_{N}^{*}(x_{0},x_{tg},t_{f})=\frac{1}{t_{f}}\left\|B^{\dagger}(v-\tilde{x})\right\|_{2}^{2},$ (20)
where $\|v\|_{\infty}\leq\overline{v}$ (16). Equation (20) thus provides the approximation to the nominal energy using the norm of the mean value $\overline{u}^{LS}$ from (15).
We now consider the malfunctioning case, which proceeds along similar lines to the nominal case. The state trajectory of the malfunctioning system (5) is given by
$x(t)=x_{0}+\int_{0}^{t}f(x(\tau))\mathrm{d}\tau+\int_{0}^{t}g_{c}(x(\tau))u_{c}(\tau)\mathrm{d}\tau+\int_{0}^{t}g_{uc}(x(\tau))u_{uc}(\tau)\mathrm{d}\tau.$ (21)
Let $B_{c}\coloneqq g_{c}(x_{0})$ and $B_{uc}\coloneqq g_{uc}(x_{0})$. Then, under finite-time reachability with $x(t_{f})=x_{tg}$,
$\tilde{x}+t_{f}B_{c}\overline{u}_{c}+t_{f}B_{uc}\overline{u}_{uc}=-\int_{0}^{t}f(x(\tau))\mathrm{d}\tau-\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t$ (22)
where $u(t)=\left[u_{c}^{\operatorname*{\mathsf{T}}}(t),u_{uc}^{\operatorname*{\mathsf{T}}}(t)\right]^{\operatorname*{\mathsf{T}}}$, and we define the mean values $\overline{u}_{c}\coloneqq\frac{1}{t_{f}}\int_{0}^{t_{f}}u_{c}(t)\mathrm{d}t$ and $\overline{u}_{uc}\coloneqq\frac{1}{t_{f}}\int_{0}^{t_{f}}u_{uc}(t)\mathrm{d}t$. Note that the right-hand side equals $v$ as defined earlier. For a fixed $v$, the above equation is linear in $\overline{u}_{c}$, and similar to the nominal case, we have the least-norm solution
$\overline{u}_{c}^{LS}=\frac{1}{t_{f}}B_{c}^{\dagger}\left(v-\tilde{x}-t_{f}B_{uc}\overline{u}_{uc}\right),$ (23)
where $\|v\|_{\infty}\leq\overline{v}$ from (16). We remind the reader that $v$ is not a fixed quantity, but the norm of $\overline{u}_{c}^{LS}$ can be used to approximate the malfunctioning energy $E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})$ in (9) as $v$ is bounded. First, we investigate conditions under which the constraint $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ is satisfied using (23) and the bound $\overline{v}$.
If $\|u_{c}(t)\|_{\infty}\leq 1$, then $\overline{u}_{c}^{LS}$ in (23) satisfies $\|\overline{u}_{c}^{LS}\|_{\infty}\leq 1$. We also note that $\|u_{uc}(t)\|_{\infty}\leq 1$ in (7) implies $\|\overline{u}_{uc}\|_{\infty}\leq 1$. As with the nominal case, reducing the expression $\|\overline{u}_{c}^{LS}\|_{\infty}\leq 1$ to a condition on $t_{f}$ is not straightforward due to the presence of $v(t_{f},u)$, as well as the nonlinear dependence of the expression on $t_{f}$. We can again check whether $\|\overline{u}_{c}^{LS}\|_{\infty}\leq 1$ for a given $t_{f}$ by using the known bound $\|v\|_{\infty}\leq\overline{v}$ in the following condition:
$\max_{\|v\|_{\infty}\leq\overline{v}}\max_{\|\overline{u}_{uc}\|_{\infty}\leq 1}\frac{1}{t_{f}}\left\|B_{c}^{\dagger}\left(v-\tilde{x}-t_{f}B_{uc}\overline{u}_{uc}\right)\right\|_{\infty}\leq 1.$ (24)
As before, if (24) is satisfied, then $\|\overline{u}_{c}^{LS}\|_{\infty}\leq 1$ is true for all $v(t_{f},u)$ and all $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$, which is necessary to ensure $u_{c}\in\operatorname*{\mathcal{U}_{c}}$. Equation (24) reduces to the following conditions:
$\max_{\|v\|_{\infty}\leq\overline{v}}\max_{\|\overline{u}_{uc}\|_{\infty}\leq 1}\frac{1}{t_{f}}\left[B_{c}^{\dagger}(v-\tilde{x}-t_{f}B_{uc}\overline{u}_{uc})\right]_{j}\leq 1$
and $\max_{\|v\|_{\infty}\leq\overline{v}}\max_{\|\overline{u}_{uc}\|_{\infty}\leq 1}-\frac{1}{t_{f}}\left[B_{c}^{\dagger}(v-\tilde{x}-t_{f}B_{uc}\overline{u}_{uc})\right]_{j}\leq 1$
for each $j=1,\ldots,m$. The objective of the inner maximum is linear in $\overline{u}_{uc}$, and hence its solution lies on the vertices of the hypercube $\|\overline{u}_{uc}\|_{\infty}\leq 1$. Following this, the objective of the outer maximum is also linear in $v$, and the solution lies on the vertices of the hypercube $\|v\|_{\infty}\leq\overline{v}$. Thus, the above conditions can be checked efficiently on the vertices of the two hypercubes $\|\overline{u}_{uc}\|_{\infty}\leq 1$ and $\|v\|_{\infty}\leq\overline{v}$ [39].
Finally, we consider the malfunctioning energy $E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})$ in (9) and note that obtaining an exact expression is intractable [37] for the general nonlinear system (5). We thus use (23) to approximate $E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})$ using Lemma 1. From (18), we have $E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})\geq t_{f}\|\overline{u}_{c}^{LS}\|_{2}^{2}$, as the mean of the optimal control signal in (9) must satisfy (23) to achieve finite-time reachability. Then, similar to the approximation for the nominal energy (19), we write the approximate malfunctioning energy
$E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})\approx t_{f}\|\overline{u}_{c}^{LS}\|_{2}^{2},$ (25)
where we recall that $\overline{u}_{c}^{LS}$ depends on $v(t_{f},u)$, which is bounded as in (16). As before, this approximation is exact when the optimal control is constant, i.e., $u_{c}^{*}(t)=\overline{u}_{c}^{LS}$ for all $t\in[0,t_{f}]$, which is true in the case of linear, driftless systems as shown in [34]. Under the assumption that this approximation is exact, we obtain
$E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})=\frac{1}{t_{f}}\left\|B_{c}^{\dagger}\left(v-\tilde{x}-t_{f}B_{uc}\overline{u}_{uc}\right)\right\|_{2}^{2},$ (26)
with $\|v\|_{\infty}\leq\overline{v}$.
Equations (20) and (26) provide the approximations for the nominal and malfunctioning energies and are obtained using Proposition 1 and Lemma 1. The conditions (17) and (24) are used to check whether the constraints $u\in\operatorname*{\mathcal{U}}$ and $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ are satisfied for a given $t_{f}$. In the next subsection, we use these results to obtain an approximate upper bound on the worst-case total energy $\overline{E}_{T}(x_{0},x_{tg},t_{f})$ defined in (11).
3.2 Worst-case Total Energy
We now consider the worst-case total energy $\overline{E}_{T}(x_{0},x_{tg},t_{f})$, defined in (11). Note that the expression satisfied by the mean of the controlled input $\overline{u}_{c}^{LS}$ in (23) and the approximate malfunctioning energy (26) depend on the uncontrolled input $u_{uc}$. The maximal effect of $u_{uc}$ is thus quantified by the worst-case total energy, representing the energy used by all actuators of the malfunctioning system. The following proposition provides an upper bound on the worst-case total energy based on the approximations (20) and (26).
Proposition 2 (Worst-case Total Energy).
Under the assumption that the approximations for the nominal (20) and malfunctioning (26) energies are exact, the worst-case total energy $\overline{E}_{T}(x_{0},x_{tg},t_{f})$ is bounded from above as follows:
$\overline{E}_{T}(x_{0},x_{tg},t_{f})\leq\frac{1}{t_{f}}\left\|B_{c}^{\dagger}(v-\tilde{x})\right\|_{2}^{2}+t_{f}\left(\sum_{i=1}^{p}\lambda_{i}\|V\|_{1}^{2}+p\right)+2\left\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right\|_{1},$ (27)
where $p$ is the dimension of the uncontrolled input $u_{uc}$, and $B_{uc}^{T}B_{uc}$ has the spectral decomposition $B_{uc}^{T}B_{uc}=V\Lambda V^{T}$, with $\Lambda=\operatorname*{diag}\{\lambda_{1},\ldots,\lambda_{p}\}$.
Proof.
From the definition of the worst-case total energy (11) and the expression for the malfunctioning energy (26),
$\overline{E}_{T}(x_{0},x_{tg},t_{f})=\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}\frac{1}{t_{f}}\left\|B_{c}^{\dagger}\left(v-\tilde{x}-t_{f}B_{uc}\overline{u}_{uc}\right)\right\|_{2}^{2}+\int_{0}^{t_{f}}u_{uc}^{T}(t)u_{uc}(t)\mathrm{d}t.$
Obtaining a closed-form analytical expression for the above problem is not straightforward, and hence we aim to bound these terms from above. Splitting the supremum over three different terms, we obtain the following bound:
$\overline{E}_{T}(x_{0},x_{tg},t_{f})\leq$
$\frac{1}{t_{f}}\left\|B_{c}^{\dagger}(v-\tilde{x})\right\|_{2}^{2}+\underbrace{\sup_{\|\overline{u}_{uc}\|_{\infty}\leq 1}t_{f}\overline{u}_{uc}^{T}B_{uc}^{T}B_{uc}\overline{u}_{uc}}_{=T_{1}}+\underbrace{\sup_{\|\overline{u}_{uc}\|_{\infty}\leq 1}2(\tilde{x}-v)^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}}_{=T_{2}}$
$+\underbrace{\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}\int_{0}^{t_{f}}u_{uc}^{T}(t)u_{uc}(t)\mathrm{d}t}_{=T_{3}},$ (28)
where $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$ implies the constraint $\|\overline{u}_{uc}\|_{\infty}\leq 1$ holds in the terms $T_{1}$ and $T_{2}$. We consider each term in (28) separately, beginning with $T_{1}$. Let the spectral decomposition of $B_{uc}^{T}B_{uc}$ be $B_{uc}^{T}B_{uc}=V\Lambda V^{T}$, where $\Lambda=\operatorname*{diag}\{\lambda_{1},\ldots,\lambda_{p}\}$ and $V$ is the orthonormal matrix of eigenvectors of $B_{uc}^{T}B_{uc}$. Next, define $q\coloneqq V^{T}\overline{u}_{uc}$. Using the sub-multiplicative property of norms, $\|q\|_{\infty}\leq\|V^{T}\|_{\infty}\|\overline{u}_{uc}\|_{\infty}\leq\|V\|_{1}$ since $\|\overline{u}_{uc}\|_{\infty}\leq 1$. Due to the use of the sub-multiplicative property, the set $\left\{\overline{u}_{uc}:\|\overline{u}_{uc}\|_{\infty}\leq 1\right\}$ is a subset of the set $\left\{q:\|q\|_{\infty}\leq\|V\|_{1}\right\}$, where $q=V^{T}\overline{u}_{uc}$. This property results in the following upper bound for $T_{1}$:
$T_{1}\leq t_{f}\sup_{\|q\|_{\infty}\leq\|V\|_{1}}q^{T}\Lambda q=t_{f}\sup_{\|q\|_{\infty}\leq\|V\|_{1}}\sum_{i=1}^{p}\lambda_{i}q_{i}^{2}=t_{f}\sum_{i=1}^{p}\lambda_{i}\|V\|_{1}^{2}.$ (29)
In (29), each component of the $q^{*}$ achieving the supremum takes on the maximum possible value, i.e., $q_{i}^{*}=\pm\|V\|_{1}$. Next, the term $T_{2}$ can be rewritten as the problem $\sup_{\|\overline{u}_{uc}\|_{\infty}\leq 1}\beta^{T}\overline{u}_{uc}=\|\beta\|_{1}$, where $\beta=2B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(\tilde{x}-v)$ and the optimal $\overline{u}_{uc}$ for this particular term is $\overline{u}_{uc}^{*}=\operatorname*{sign}(\beta)=\operatorname*{sign}\left(2B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(\tilde{x}-v)\right)$. Then,
$T_{2}=2\left\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right\|_{1}.$ (30)
Finally, $T_{3}$ is clearly maximized by each component of the uncontrolled input $u_{uc}$ taking on the maximal value of $+1$ or $-1$ at each time instant $t$, so that $u_{uc}^{T}(t)u_{uc}(t)=p$ for all $t$, and
$T_{3}=t_{f}p.$ (31)
Using (29), (30) and (31) in (28), the result (27) follows. ∎
Proposition 2 thus provides an upper bound on the worst-case total energy, assuming the approximations for the nominal (20) and malfunctioning (26) energies are exact. In the next subsection, we consider the special case when control authority is lost over one actuator, i.e., when $p=1$. In this particular case, we derive a closed-form expression instead of an upper bound for $\overline{E}_{T}(x_{0},x_{tg},t_{f})$, assuming the above approximations are still exact. We subsequently use this closed-form expression to obtain a bound on the energetic resilience metric (12) in Section 4.
3.3 Loss of Authority over One Actuator
When control authority is lost over one actuator, we have $p=1$, $u_{uc}(t)\in\mathbb{R}$ and $B_{uc}\in\mathbb{R}^{n}$. Then, the worst-case total energy $\overline{E}_{T}(x_{0},x_{tg},t_{f})$ is given by the following proposition.
Proposition 3.
Under loss of control authority over one actuator, i.e., when $p=1$, and under the assumption that the approximations for the nominal (20) and malfunctioning (26) energies are exact, the worst-case total energy $\overline{E}_{T}(x_{0},x_{tg},t_{f})$ is expressed as
$\overline{E}_{T}(x_{0},x_{tg},t_{f})=\frac{1}{t_{f}}\left\|B_{c}^{\dagger}(v-\tilde{x})\right\|_{2}^{2}+t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right)+2\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right|.$ (32)
Proof.
From the definition (11) and the expression for the malfunctioning energy (26),
$\overline{E}_{T}(x_{0},x_{tg},t_{f})=\frac{1}{t_{f}}\left\|B_{c}^{\dagger}(v-\tilde{x})\right\|_{2}^{2}$ $+\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}\biggl{\{}t_{f}B_{uc}^{T}B_{uc}\overline{u}_{uc}^{2}+2(v-\tilde{x})^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}$
$+\int_{0}^{t_{f}}u_{uc}^{2}(t)\mathrm{d}t\biggr{\}}$ (33)
The constraint $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$ is equivalent to the constraints $-1\leq\overline{u}_{uc}\leq 1$ on $\overline{u}_{uc}$. Then, the second term inside the supremum reduces to
$\sup_{-1\leq\overline{u}_{uc}\leq 1}2(v-\tilde{x})^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}=2\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right|,$
with the optimal $\overline{u}_{uc}^{*}$ for this particular term given by $\overline{u}_{uc}^{*}=\operatorname*{sign}\left(B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right)$. When $\operatorname*{sign}\left(B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right)\in\{-1,+1\}$, this choice of $\overline{u}_{uc}^{*}$ also maximizes the first term inside the supremum in (33), with $\sup_{-1\leq\overline{u}_{uc}\leq 1}t_{f}B_{uc}^{T}B_{uc}\overline{u}_{uc}^{2}=t_{f}\|B_{uc}\|_{2}^{2}$. Further, $\overline{u}_{uc}^{*}\in\{-1,+1\}$ implies that the uncontrolled input $u_{uc}^{*}(t)$ also takes on a constant value of $-1$ or $+1$ at each instant $t\in[0,t_{f}]$. This value for $u_{uc}^{*}$ maximizes the third term inside the supremum in (33), with $\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}\int_{0}^{t_{f}}u_{uc}^{2}(t)\mathrm{d}t=t_{f}$.
On the other hand, if $B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})=0$, the second term inside the supremum in (33) vanishes, leading to a smaller value for $\overline{E}_{T}(x_{0},x_{tg},t_{f})$ than if this term did not vanish. The worst-case total energy is thus achieved only when $B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\neq 0$. The uncontrolled input $u_{uc}^{*}$ maximizing each of the three terms inside the supremum in (33) is the same, and is a constant equal to $-1$ or $+1$ at almost every instant $t\in[0,t_{f}]$. Combining the results of each term above, we obtain the result (32). ∎
Remark 2.
We remind the reader that the primary results of this section rely on approximations made on the basis of Lemma 1. Further, the expressions for the different energies in (20), (26) and (27) depend on $v\equiv v(t_{f},u)$, with $\|v\|_{\infty}\leq\overline{v}$. These expressions play a key role in determining bounds on the energetic resilience metric (12), as shown in the next section.
4 Bounding the Energetic Resilience
In this section, we quantify the maximal additional energy used by the malfunctioning system (5), i.e., addressing Problem 2, for the particular case of loss of control authority over one actuator. Namely, we derive a bound on the energetic resilience metric (12), assuming the approximation for the nominal (20) and worst-case total (32) energies are exact. We note that the more general case where authority is lost over $p>1$ actuators can be addressed using (27). However, this bound is more conservative as (27) provides only an upper bound on the approximate worst-case total energy.
First, given $\tilde{x}=x_{0}-x_{tg}$, we rewrite the definition (12) as
$r_{A}(t_{f},R)$ $=\sup_{\|\tilde{x}\|_{2}\leq R}\overline{E}_{T}(x_{0},x_{tg},t_{f})-E_{N}^{*}(x_{0},x_{tg},t_{f}).$
Next, note that the nominal (20) and worst-case total (27) energies depend only on $v-\tilde{x}$, and not directly on $\tilde{x}$ itself. Then, we can use the bounds $\|\tilde{x}\|_{2}\leq R$ and $\|v\|_{\infty}\leq\overline{v}$ to obtain
$\|v-\tilde{x}\|_{2}\leq\|v\|_{2}+\|\tilde{x}\|_{2}\leq\sqrt{n}\|v\|_{\infty}+\|\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n},$
where we note that the norm inequality $\|v\|_{2}\leq\sqrt{n}\|v\|_{\infty}$ holds for $v\in\mathbb{R}^{n}$ [31, Chapter 5]. We then replace the constraint $\|\tilde{x}\|_{2}\leq R$ with $\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}$. Due to the norm inequality above, the set $\{d:\|\tilde{x}\|_{2}\leq R\}$ is a subset of the set $\{d:\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}$ where $\|v\|_{\infty}\leq\overline{v}$. This property results in the following upper bound on $r_{A}(t_{f},R)$:
$r_{A}(t_{f},R)\leq\sup_{\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}}\overline{E}_{T}(x_{0},x_{tg},t_{f})-E_{N}^{*}(x_{0},x_{tg},t_{f}).$ (34)
We can then use (34) to obtain the following bound on the resilience metric when $p=1$.
Proposition 4.
When $p=1$, the energetic resilience $r_{A}(t_{f},R)$ is bounded from above as follows:
$r_{A}(t_{f},R)\leq\frac{(R+\overline{v}\sqrt{n})^{2}}{t_{f}}\lambda_{\max}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right)$ $+2(R+\overline{v}\sqrt{n})\left\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\right\|_{2}$
$+t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right).$ (35)
Proof.
Using (20) and (32) in (34),
$r_{A}(t_{f},R)\leq t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right)$ $+2(R+\overline{v}\sqrt{n})\left\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\right\|_{2}$
$+\sup_{\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}}\frac{1}{t_{f}}(v-\tilde{x})^{T}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right)(v-\tilde{x}),$
where the second term above follows from applying the Cauchy-Schwarz inequality to bound $2\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right|$. Since $B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}$ is symmetric, we can use the Rayleigh inequality [31] and the constraint $\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}$ to bound the third term as
$\sup_{\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}}\frac{1}{t_{f}}(v-\tilde{x})^{T}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right)(v-\tilde{x})\leq\frac{(R+\overline{v}\sqrt{n})^{2}}{t_{f}}\lambda_{\max}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right),$
from which the result (35) follows. ∎
The bound in (35) thus quantifies the maximal additional energy needed for a nonlinear system which loses authority over one actuator, where we remind the reader of the approximations for the nominal (20) and worst-case total (32) energies. In the following section, we present an example that demonstrates the use of this metric in characterizing the increased energy required without significant conservatism.
5 Simulation Example
We now present an example of a nonlinear system demonstrating the accuracy of the bound (35). In particular, we show that this bound is useful in characterizing the maximal additional energy used by a system (5) to achieve a target state, while compensating for the effect of uncontrolled inputs $u_{uc}$.
We consider the following system which contains a number of representative nonlinear functions that commonly occur in practical systems:
$\begin{split}\dot{x}_{1}&=x_{2}+\sin(x_{1})+(1+\sin^{2}(x_{2}))u_{1}+3u_{2},\\ \dot{x}_{2}&=x_{1}-x_{2}-\sin(2x_{2})+2u_{1}+(1-2\cos(x_{1}))u_{2}.\end{split}$ (36)
This system can be written in the form of equation (1), with
$f(x)=\begin{bmatrix}x_{2}+\sin(x_{1})\\ x_{1}-x_{2}-\sin(2x_{2})\end{bmatrix};\hskip 10.00002ptg(x)=\begin{bmatrix}(1+\sin^{2}(x_{2}))&3\\ 2&(1-2\cos(x_{1}))\end{bmatrix}.$
We now find constants $D_{f}$ and $D_{g}$ such that (2) and (3) are satisfied. First, $f(x)$ can be written as
$f(x)=\begin{bmatrix}0&\phantom{-}1\\ 1&-1\end{bmatrix}x+\begin{bmatrix}\sin(x_{1})\\ -\sin(2x_{2})\end{bmatrix}.$ (37)
Note that the Lipschitz constant of $f(x)$ in the $\infty$-norm satisfies (2) for the constant $D_{f}$. The Lipschitz constant of the linear term in (37) in the $\infty$-norm is $\left\|\begin{bmatrix}0&\phantom{-}1\\ 1&-1\end{bmatrix}\right\|_{\infty}=2$. Next, the Lipschitz constant of $\sin(x_{1})$ is $1$ and the Lipschitz constant of $-\sin(2x_{2})$ is $2$. Then, the Lipschitz constant of $\begin{bmatrix}\sin(x_{1})\\ -\sin(2x_{2})\end{bmatrix}$ in the $\infty$-norm is $\max\{1,2\}=2$. Combining the linear and nonlinear terms, the Lipschitz constant of $f(x)$ above is $42+\max\{1,2\}=$ 4, from which $D_{f}=4$ satisfies (2).
Figure 1: Difference between nominal and malfunctioning energies, compared to the upper bound on the metric $r_{A}(t_{f},R)$. The thinner dashed lines plot $E_{T}-E_{N}$ for a variety of uncontrolled inputs $u_{uc}$.
Next, we find the constant $D_{g}$ by noting that
$g(x)=\begin{bmatrix}1&3\\ 2&1\end{bmatrix}+\begin{bmatrix}\sin^{2}(x_{2})&0\\ 0&-2\cos(x_{1})\end{bmatrix}=g_{0}+d_{g}(x).$
It is clear that $\|d_{g}(x)\|_{\infty}\leq 2$ for all $x$, and thus $D_{g}=2$.
In this example, we assume that a malfunction causes the system to lose authority over $u_{2}$, so that $u_{c}=u_{1}$ and $u_{uc}=u_{2}$, with $g(x)$ split according to (5). We fix $t_{f}=1$ and vary the distance $R$ of the initial condition from the target state, which is chosen to be $x_{tg}=[0,0]^{T}$. For notational simplicity, we denote $E_{N}\equiv E_{N}^{*}(x_{0},x_{tg},t_{f})$ for the approximate nominal energy, $E_{T}\equiv E_{T}(x_{0},x_{tg},t_{f},u_{uc})$ for the approximate total energy, and $\overline{E}_{T}\equiv\overline{E}_{T}(x_{0},x_{tg},t_{f})$ for the approximate worst-case total energy.
Fig. 1 compares the difference between the total energies and the nominal energy, to the upper bound on the metric $r_{A}(t_{f},R)$ derived in (35). This upper bound is shown as the thick red line, while the worst-case difference in energies $\overline{E}_{T}-E_{N}$ using (20) and (32) is shown as the thick blue line. It is clear that even in the worst-case, this difference in energies is upper bounded by (35). Further, the bound is meaningfully accurate, with a relatively small error. In particular, the relative error between $\overline{E}_{T}-E_{N}$ and the bound on $r_{A}(t_{f},R)$ is consistently between just $3\%$ and $4\%$ for the range of $R$ considered.
In Fig. 1, we also plot the difference between the total and nominal energies, $E_{T}-E_{N}$ for a large number of uncontrolled inputs $u_{uc}$. This includes sinusoidal inputs, constant inputs of different amplitudes, as well as decaying exponentials and sinusoids. These are shown as the thinner dashed lines. As expected, this difference is smaller than both the worst-case difference $\overline{E}_{T}-E_{N}$ and the upper bound on the metric. Once again, the worst-case difference and the upper bound on $r_{A}(t_{f},R)$ are meaningful in characterizing the additional energy required to achieve finite-time reachability under a partial loss of control authority.
We note that for $R=5$, the upper bound on $r_{A}(t_{f},R)$ is approximately $2640$ units of energy. For the different malfunctioning inputs chosen here, $E_{T}-E_{N}$ ranges from approximately from $2240$ to $2420$ units of energy. Further, the worst-case difference in energies $\overline{E}_{T}-E_{N}$ is approximately $2540$ units. Note that the upper bound on $r_{A}(t_{f},R)$ is only $3.9\%$ larger than $\overline{E}_{T}-E_{N}$. Further, $\overline{E}_{T}-E_{N}$ is at most $13\%$ larger than the actual difference in energies $E_{T}-E_{N}$ for particular choices of the uncontrolled input $u_{uc}$. Thus, $r_{A}(t_{f},R)$ is a useful quantity to approximate the increased energy required without significant conservatism, despite our knowledge of only its upper bound.
6 Conclusions
In this article, we introduce a metric to bound the maximal additional energy used by a nonlinear dynamical system in achieving a finite-time reachability task, when it loses authority over a subset of its actuators. This maximal energy is compared to the nominal energy expended when no authority is lost over the inputs. We first derive an expression satisfied by the mean of the control signal used in the nominal and malfunctioning case. We then show that the energy in this mean value can be used to approximate the energy in the control signal in both cases. Under this approximation, we derive a bound on the worst-case malfunctioning energy over all possible malfunctioning inputs, and obtain an exact expression in the case when authority is lost over only one actuator. This exact expression is used to derive a bound on the maximal additional energy when authority is lost over one actuator. A simulation example on a nonlinear dynamical system demonstrates the applicability of this metric in characterizing the increased energy required, despite the approximations used for the control energies.
Our future objectives include developing a framework for quantifying energetic resilience of nonlinear systems without using approximations for nominal and malfunctioning energies. This may be possible to accomplish through the use of techniques such as control-Lyapunov functions or considering controllers with a feedback linearization structure [35].
Appendix A Proof of Proposition 1
Recall that
$v$ $=-\int_{0}^{t_{f}}f(x(t))\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t$
$=-t_{f}f_{0}-\int_{0}^{t_{f}}d_{f}(x(t))\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t$
using (2). Then,
$\|v\|_{\infty}\leq t_{f}\|f_{0}\|_{\infty}+\left\|\int_{0}^{t_{f}}d_{f}(x(t))\mathrm{d}t\right\|_{\infty}+\left\|\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t\right\|_{\infty}.$
Using Jensen’s inequality to move the norm inside the integral, and subsequently using conditions (2) and (3), we note that
$\left\|\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t\right\|_{\infty}\leq\int_{0}^{t_{f}}\left\|d_{g}(x(t))u(t)\right\|_{\infty}\mathrm{d}t\leq t_{f}D_{g},$
and $\left\|\int_{0}^{t_{f}}d_{f}(x(t))\mathrm{d}t\right\|_{\infty}\leq\int_{0}^{t_{f}}\|d_{f}(x(t))\|_{\infty}\mathrm{d}t\leq D_{f}\int_{0}^{t_{f}}\|x(t)-x_{0}\|_{\infty},$
where we use the sub-multiplicative property of matrix norms and the input constraint $\|u(t)\|_{\infty}\leq 1$ for all $t$ (4). Thus,
$\|v\|_{\infty}\leq t_{f}(\|f_{0}\|_{\infty}+D_{g})+D_{f}\int_{0}^{t_{f}}\|x(t)-x_{0}\|_{\infty}\mathrm{d}t.$ (38)
We require a method to evaluate the third term in (38). From (13),
$x(t)-x_{0}=tf_{0}+\int_{0}^{t}d_{f}(x(\tau))\mathrm{d}\tau+t_{f}B\overline{u}+\int_{0}^{t}d_{g}(x(\tau))u(\tau)\mathrm{d}\tau.$
Applying $\|\cdot\|_{\infty}$ on both sides and following steps similar to those used in obtaining (38), we obtain
$\|x(t)-x_{0}\|_{\infty}\leq t(\|f_{0}\|_{\infty}+\|B\|_{\infty}+D_{g})+D_{f}\int_{0}^{t}\|x(\tau)-x_{0}\|_{\infty}\mathrm{d}\tau.$ (39)
Define $Q(t)\coloneqq\int_{0}^{t}\|x(\tau)-x_{0}\|_{\infty}\mathrm{d}\tau$ and $c\coloneqq\|f_{0}\|_{\infty}+\|B\|_{\infty}+D_{g}$. Note that $Q(0)=0$. In (38), we require $Q(t_{f})$ in the third term. From the first fundamental theorem of calculus [40, Theorem 5.1], $\dot{Q}(t)=\|x(t)-x_{0}\|_{\infty}$. Then, (39) reduces to
$\dot{Q}(t)\leq ct+D_{f}Q(t),~{}Q(0)=0.$ (40)
Let $Q_{1}(t)\coloneqq ct+D_{f}Q(t),Q_{1}(0)=0$ so that $\dot{Q}(t)\leq Q_{1}(t)$. Then, $\dot{Q}_{1}(t)=c+D_{f}\dot{Q}(t)\leq c+D_{f}Q_{1}(t)$. Next, let $Q_{2}(t)\coloneqq c+D_{f}Q_{1}(t),Q_{2}(0)=c$ so that $\dot{Q}_{1}(t)\leq Q_{2}(t)$. Then, $\dot{Q}_{2}(t)=D_{f}\dot{Q}_{1}(t)\leq D_{f}Q_{2}(t)$. We thus have the differential inequality
$\dot{Q}_{2}(t)\leq D_{f}Q_{2}(t),~{}Q_{2}(0)=c.$
We now invoke Grönwall’s inequality [33, Theorem 1.2.1], using which we obtain
$Q_{2}(t)\leq Q_{2}(0)e^{D_{f}t}=ce^{D_{f}t}.$
Substituting back in $Q_{1}(t)$ and $Q(t)$, we have
$Q_{1}(t)$ $\leq\frac{c(e^{D_{f}t}-1)}{D_{f}}$
and $\text{and }Q(t)$ $\leq\frac{c(e^{D_{f}t}-1-D_{f}t)}{D_{f}^{2}}.$
Thus,
$Q(t_{f})\leq\frac{c(e^{t_{f}D_{f}}-1-t_{f}D_{f})}{D_{f}^{2}}.$ (41)
Substituting the bound (41) on $Q(t_{f})=\int_{0}^{t_{f}}\|x(t)-x_{0}\|_{\infty}\mathrm{d}t$ in the third term of (38),
$\|v\|_{\infty}\leq t_{f}(\|f_{0}\|_{\infty}+D_{g})+\frac{c}{D_{f}}(e^{t_{f}D_{f}}-1-t_{f}D_{f}),$
from which the result (16) of Proposition 1 follows using the definition of $c$. $\blacksquare$